OpenStudy (anonymous):

A package of four plastic hats cost $8. What is the unit price of a hat? Write your answer as a comparison statement. Ill give medal + fan

OpenStudy (anonymous):

ill give medal + fan soo plzzz help :)

OpenStudy (emmamink):

2 dollars

OpenStudy (emmamink):

how I got the answer: take 8 dollars and divide it by the number of hats. 8 divided by 4 is 2!

OpenStudy (texaschic101):

comparison statement ? would that be the same as a proportion ?

OpenStudy (emmamink):

So I'm waiting for a medal and to be fanned!

OpenStudy (anonymous):

i guess texaschcick

OpenStudy (texaschic101):

4/8 = 1/x --- (4 hats for $8 = 1 hat for x $) cross multiply 4x = 8 x = 2 Not sure if this is a comparison

OpenStudy (anonymous):

ok